OpenStudy (anonymous):

tan^4 h -1 I need to express this in terms of sin h

OpenStudy (anonymous):

\[\tan ^{4}(h-1)=\sin ^{4}(h-1)\div \cos ^{4}(h-1)=\sin ^{4}(h-1)\div (1-\sin ^{2}(h-1))^{2}\] is it sufficient or you need to simplify more...?

OpenStudy (anonymous):

or is it \[tan^4(h)-1\]?

OpenStudy (anonymous):

I don't know if I need to simplify more. Can you? and its (h-1)